# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ II.2 Mutants in both or either subpopulations
Few remarks are in order. ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants. In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in sum-definition. In the light of above remarks, we now present the following set of equivalent definitions (aptly termed or-definition) where one of $\epsilon_{x}$ or $\epsilon_{y}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
@@
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{or}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
@@
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both the equalities in (i) hold,
    $\text{if}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{then}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}~{}\text{then}~{}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (13)
@@
    $\text{either}~{}~{}~{}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
@@
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998, 2003) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
@@
We notice that (6) directly matches with equilibrium conditions in Definition 3b. So obviously, ESS under or-definition implies NE; and strict NE trivially implies ESS. However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’. However, this apparent anomaly is prevented since the definition allows for one of $\epsilon_{x}$ or $\epsilon_{y}$ to be zero independently. Specifically, in the case pointed out, mathematically setting $\epsilon_{x}=0$ (and $\epsilon_{y}\neq 0$), physically means that no mutants in X-subpopulation are allowed, i.e., no ${\bf x}\neq{\bf{\hat{x}}}$ can be used to satisfy $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, and hence the first inequality in (11) for infinitesimally small $\epsilon_{y}$. Owing to the conjunction ‘or’, the second inequality in (11) must be checked which anyway can’t be satisfied as we have assumed that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. In conclusion, allowing $\epsilon_{x}$ and $\epsilon_{y}$ to vanish (but not simultaneously) prevents states that are not NE’s from qualifying as ESS’s. Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
@@
Consider Definition 3a version of or-definition. If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$. Consequently, setting any of the $\epsilon_{x}$ or $\epsilon_{y}$ to zero is not permissible, as it renders both conditions meaningless. For instance, if we set $\epsilon_{y}=0$, then both the comparisons $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid because $(\epsilon_{x},0)$ physically means there is no mutant (no choice of ${\bf y})$ in the Y-population. Thus, if you would want to allow mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS mutants have to be present in both X- and Y-subpopulations—this perspective may appositely be called the simultaneous deviation perspective. This view is adopted while defining the Nash–Pareto pair in Hofbuaer and Sigmund Hofbauer and Sigmund (1998). The alternate perspective—which we term face-value perspective is to permit all deviations outlined explicitly by the or-definition [as specified in Definition 3a, $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE. This view was adopted by Cressman Cressman and Tao (2014); Cressman (1992) while defining 2ESS. It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
@@ II.3 Equivalent definitions
Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all infinitesimal $(\epsilon_{x},\epsilon_{y})\neq(0,0)$, inequalities in (11) are valid. First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there always is some permissible value of $(\epsilon_{x},\epsilon_{y})$ for which the inequalities in (11) do not hold. For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $\epsilon_{x}\neq 0$ and $\epsilon_{y}=0$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid because there is no mutant in Y-population hence the comparison in the second condition in (11) won’t make any physical sense. Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—across the allowable ranges of $\epsilon_{x}$ and $\epsilon_{y}$ to satisfy the inequalities in (11) is:
$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
Thus, four explicit possibilities arise from (15) as detailed below. We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition in (12)) must be satisfied with the natural requirement of conditions in (13) in the relevant case.
@@
    One can trivially see that either condition— $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$—is true when (11) holds good, or
@@
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
@@
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
@@
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$, straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the conditions in (13) must follow.
Above argument establishes that Definition 3a implies Definition 3b. Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent. We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESS’s possible are the ones that are strict NE.
Next let us show that Definitions 3a and 3c imply each other. Starting with Eq. (11), we multiply both sides of the first inequality of Eq. (11) by $\epsilon_{x}$ and add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides. Similarly, we multiply both sides of the second inequality of Eq. (11) by $\epsilon_{y}$ and add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides. The two resultant inequality expressions can be rearranged to read:

### v3

@@ II.2 Mutants in both or either subpopulations
Few remarks are in order. ESSes defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and demanding only that residents outperform the mutants in at least one subpopulation. In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in sum-definition. In the light of above remarks, we now present the following set of equivalent definitions (aptly termed or-definition) where one of $\epsilon_{x}$ or $\epsilon_{y}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
@@
    ${\color[rgb]{0,0,0}\epsilon_{x}}\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>{\color[rgb]{0,0,0}\epsilon_{x}}\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\color[rgb]{0,0,0}\epsilon_{y}}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>{\color[rgb]{0,0,0}\epsilon_{y}}\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
    This definition is provided in Ref. Cressman (1996) in the context of asymmetric games.
@@
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both inequalities in (i) hold as equalities, following condition holds
    ${\color[rgb]{0,0,0}\text{either}\leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.}$ (13)
@@
    $\text{either}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
@@
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS in Definition 3c is known as two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature. Some readers may be reminded of the Nash–Pareto pair while reading Definition 3b; however, there is a subtle difference between the two (see Appendix A.2).
@@
We notice that (6) directly matches with equilibrium conditions in Definition 3b. So obviously, ESS under or-definition implies NE; and strict NE trivially implies ESS. However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’. However, this apparent anomaly is prevented since the definition allows all the mutant state $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$. For instance, if we consider a mutant state of the form $(\mathbf{\hat{x}},\mathbf{y})$, which is distinct from $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the first inequality in (11) cannot hold. Moreover, in this case the second inequality simplifies to $\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}>\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ preventing the erroneous conclusion. In conclusion, allowing all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ prevents states that are not NEs from qualifying as ESSes. Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
@@
Consider Definition 3a version of or-definition. If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$. Consequently, setting the mutant state $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$ should not be permissible, as it renders both conditions meaningless. For instance, if we take $(\mathbf{x},\mathbf{\hat{y}})$ as mutant, then both the inequalities $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid (as the first one contradicts $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and the second one trivially becomes an equality). Thus, if one would want to allow for a mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS one cannot allow for mutants of the form $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$. One can also argue that allowing for mixed 2ESS needs the restriction $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$. This is because the simplified condition after putting the mixed Nash equilibrium criteria in (11) becomes: $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$, which can never be satisfied if one allows for either of the $\epsilon_{x}$ or $\epsilon_{y}$ to go to zero. This perspective (ad hoc simultaneous imposition of $\mathbf{x}\neq\mathbf{\hat{x}},\leavevmode\nobreak\ \mathbf{y}\neq\mathbf{\hat{y}},\leavevmode\nobreak\ \epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$) may appositely be called the simultaneous deviation perspective. This view is adopted while defining the Nash–Pareto pair in Hofbauer and Sigmund Hofbauer and Sigmund (1998).
An alternative perspective—which we term face-value perspective is to permit all mutants (including $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$) as outlined explicitly by the or-definition [as specified in Definition 3a, with $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$$0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE. This view was adopted by Cressman Cressman and Tao (2014); Cressman (1992) while defining 2ESS. It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
We would like to re-emphasize that by ‘face-value perspective’, we mean strictly following the definitions without imposing any additional constraints (whether they involve allowing mutant states like $(\mathbf{\hat{x}},\mathbf{y})$ or letting $\epsilon_{y}$ go to zero). On the other hand, adoption of ‘simultaneous deviations perspective’, where one puts additional constraint on the allowed mutant states and their fractions, allows for the possibility of existence of mixed ESS in bimatrix games. At this point, one might question, why we are emphasizing so much on the existence of a mixed ESS in bimatrix games, given that it can never be asymptotically stable under replicator dynamics (see Section III.3). The reason is that stability depends on the choice of evolutionary dynamics. As discussed in Section III.3, while a mixed ESS is not stable under replicator dynamics, it can be stable under other evolutionary dynamics. Since replicator dynamics is not the only possible evolutionary process, this opens the possibility that certain mixed states—though unstable under replicator dynamics—can still qualify as an ESS and exhibit dynamical stability under alternative dynamics.
@@ II.3 Equivalent definitions
Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all mutants $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, inequalities in (11) are valid. First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}<\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there is always some mutant $(\mathbf{x},\mathbf{y})$ for which the inequalities in (11) do not hold. For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $(\mathbf{x},\mathbf{\hat{y}})$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid. Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—for all allowable mutants $(\mathbf{x},\mathbf{y})$ to satisfy the inequalities in (11) is:
$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
Thus, four explicit possibilities arise from (15) as detailed below. We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition (12)) must be satisfied with the natural requirement of condition (13) in the relevant case. One can trivially see that following is true when when (11) holds good:
@@
    either condition—$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
@@
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
@@
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
@@
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$, $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the condition (13) must follow.
Above argument establishes that Definition 3a implies Definition 3b. Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be an ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent. We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESSes possible are the ones that are strict NE.
Next let us show that Definitions 3a and 3c imply each other. We start by assuming that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ satisfies Definition 3a. We note every $(\mathbf{x}_{1},\mathbf{y}_{1})$ close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can written as $(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}},\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ where $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small. Let us first show that $\epsilon_{x}$ and $\epsilon_{y}$ can be chosen independent of $(\mathbf{x},\mathbf{y})$ in a continuous manner. Consider a set, say, $\Gamma$, which is the union of all faces of $\Sigma_{m}\times\Sigma_{n}$ that does not contain $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. We note that $\Gamma\equiv\Gamma_{x}\times\Gamma_{y}$, where $\Gamma_{x}\equiv\{\mathbf{x}\in\Sigma_{m}:x_{i}=0\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ i\in{\rm supp}(\mathbf{\hat{x}})\}$ and $\Gamma_{y}\equiv\{\mathbf{y}\in\Sigma_{n}:y_{i}=0\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ i\in{\rm supp}(\mathbf{\hat{y}})\}$, is compact Munkres (2000); Hofbauer and Sigmund (1998). By definition, for every $(\mathbf{x},\mathbf{y})\in\Gamma$, (11) holds for all $(\epsilon_{x},\epsilon_{y})$ such that $\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$. We can construct a continuous function Cressman (1996), $\epsilon_{(\mathbf{x},\mathbf{y})}\in(0,1]$ and, since the set $\Gamma$ is compact, $\epsilon_{\min}\equiv\min\{\epsilon_{(\mathbf{x},\mathbf{y})}:(\mathbf{x},\mathbf{y})\in\Gamma\}$ is strictly positive. Therefore (11) holds for all $\epsilon_{x},\epsilon_{y}<\epsilon_{\min}$ but $(\epsilon_{x},\epsilon_{y})\neq(0,0)$. We can now do the following steps to prove the equivalence: Starting with (11), we add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides of the first inequality of (11). Similarly, we add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides of the second inequality of (11). The two resultant inequality expressions can be rearranged to read: